HUBBLE SPACE TELESCOPE DAILY REPORT #1344 PERIOD COVERED: 0000Z (UTC) 03/09/95 - 0000Z (UTC) 03/10/95 Daily Status Report as of 069/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 5353 (Low-Mass Binaries in the Hyades: Cycle 4 High) The WF/PC-2 was used to make images of the low-mass, late M-Class dwarf star RHY060 which is located in the Hyades cluster. Some of its companion stars may be brown dwarfs. The observations were completed as scheduled, and no problems were reported. 1.2 Completed FOS/BL 5658 (White Dwarf Calibration) The Faint Object Spectrograph (Blue Side) was used to make absolute sensitivity calibrations of the detectors using standard white dwarf star GD71. The observations were executed as planned, and no anomalies were noted. 1.3 Completed Two Sets of WF/PC-2 6140 (WF/PC-2 Cycle 4 Calibration: Ramp filters, Internal Calibration) The WF/PC-2 was used to make visual flat calibration images as part of a ramp filter calibration test. The observations were completed as planned and no problems were reported. 1.4 Completed WF/PC-2 5302 (The Size and Morphology of Normal Galaxies at Z ~ 3) The WF/PC-2 was used to make Lyman_Alpha emission observations of galaxy PKS0528-25 which is known to have a redshift of Z ~ 3. It is hoped that the nature of this source can be determined. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed FOS/BL 5444 (A UV Spectral Library of Metal- Poor Massive Stars) The FOS (Blue Side) was used for a snapshot program to obtain ultraviolet spectra of AV469 a metal-poor star in the LMC. It will be useful for stellar atmosphere and stellar evolution studies. The observations completed without incident. 1.6 Completed WF/PC-2 5655 (WF/PC-2 Cycle 4 Calibration: Internal Monitor. 2: Flats Part 2) The WF/PC-2 was used to make a visual flat. The observation was completed as planned, and no problems were reported. 1.7 Completed WF/PC-2 6145 (Ultraviolet Imagery of Jovian Aurora to support FUV Spectroscopy) The WF-PC-2 was used to make ultraviolet images of the Jovian aurora on the planet Jupiter. Concurrent observations were made in the far ultraviolet by the Astro-2 Shuttle mission with the HUT camera. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed Two Sets of WF/PC-2 5562 (WF/PC-2 Cycle 4 Calibration: Internal Monitor 3: Darks) The WF/PC-2 was used to obtain several 30-minute long, dark exposures to provide ongoing calibration of the CCD dark current and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed WF/PC-2 5742 (Globular Cluster Cores, WF/PC-2 GTO Augmentation, Cycle 4) The WF/PC-2 was used to obtain images of the core of globular cluster NGC-6341 to look for evidence of collapsed cusps. Offset images were also taken. The observations were completed as planned, and no problems were reported. 1.10 Completed Two Sets of GHRS/2 5671 (GHRS Dark Noise Monitor: Cycle 4) The High Resolution Spectrograph (Side Two) was used to make dark calibration images. These were strictly internal tests for routine monitoring of detector dark noise. The observations were executed as scheduled, and no anomalies were noted. 1.11 Completed WF/PC-2 5479 (Subarcsecond Structure in Nearby AGNs) The WF/PC-2 was used to make snapshot observations of the nearby active galactic nuclei of MARK290 to study the interactions between the central engine, the stars, and the interstellar matter in the AGN. The observations were executed as scheduled, and no anomalies were noted. 1.12 Completed FOC/96 5409 (Direct UV and Spectral Tomogrpah Imaging of Alpha Orionis, Cycle 4 High) The Faint Object Camera (f/96) was used to make ultraviolet observations of the super giant star Alpha Orionis (aka, HD39801, Betelgeuse). The observations were executed as scheduled, and no anomalies were noted. 1.13 Completed WF/PC-2 5370 (HST Medium-Deep Survey: Cycle 4 Medium, High Galactic Latitude) The WF/PC-2 was used in parallel with the prime science observations to make high galactic latitude observations as part of a medium-deep sky survey. The observations were completed as planned, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 8 Successful : 8 Scheduled Reacquisitions : 8 Successful : 8 2.2 FHST Updates: Scheduled : 20 Successful : 20 2.3 Significant HSTARs: None 2.4 Operations Notes: Service Reconfiguration was executed for the SADE-1R at 068/1925Z. The system was configured to the "P" format for the first SA slew at 069/1117Z. A 90 to 60 degree SA slew was executed with SADE-1R taking a duration of 5 minutes and 57 seconds at 069/1144Z. Nominal operations were encountered throughout the slew. SADE-2 has been powered down. 3.0 ANOMALY STATUS: 3.1 Significant New Anomalies: None 3.2 Significant Unresolved Anomalies: None 4.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-